6550 400 5500 15100 9. Genetics Experiment In one of Mendels experiments with plants, 1064 ofspring con- sisted of 787 plants with long stems. According to Mendels theory, 3/4 of the offspring plants should have long stems. Assuming that Mendels proportion of 3/4 is correct, find the probability of getting 787or fewer plants with long stems among 1064 offspring plants. Based on the result, is 787 offspring plants with long stems unusually low? What does the result imply about Mendels claimed proportion of 3/4? k There is an 80% chance that a prospective

How would I do this problem? Iknow n=1064, and 3/4=.75 but how do I set up the rest?
Answer #1

Solution:-

The probability of getting 787 or fewer plants with long stems among 1064 plants is 0.218.

P = 0.75

T 787 n 1064

\hat{p}= 0.73966

z= \frac{\hat{p}-p}{\sqrt{\frac{p(1-p)}{n}}}

z= \frac{0.73966-0.75}{\sqrt{\frac{0.75(1-0.75)}{1064}}}

z = - 0.779

P( z < - 0.779) = 0.218

Since the p-value is greater than 0.75, hence we have sufficient evidence in the favor of Mandel's claimed proportion of 3/4.